OpenStudy (anonymous):

When a certain number is divided by 8 there is no remainder. If when the number is divided by 16 there is a remainder, the remainder must be?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

The remainder must be 8.

OpenStudy (anonymous):

i would be a miracle to divide a number by 16 and get a remainder of 24!

OpenStudy (anonymous):

but 24 is a good example of a number divisible by 8 and not by 16, so you can use it as an example

OpenStudy (mathmale):

While there may be more sophisticated ways in which to approach this problem, you might want to experiment with actual numbers to determine whether or not some pattern emerges. Suppose the "certain number," the dividend, is x. Then x/8 results in no remainder. Just supposing that x were 64, then x/8 = 64/8 = 8 with no remainder. Dividing x (or 64) by 16 produces no remainder. Thus, the "certain number" is not 64. satellite suggests using as the "certain number" 24. Do that. Go through the same reasoning as illustrated above. What is the remainder when you divide 16 into 24?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

thankx guize....